You are here: Foswiki>ARGOS Web>Commissioning>DiaryRun5>NightTime20150226 (18 Mar 2015, TommasoMazzoni)Edit Attach

Daytime activity

DARK LGSW: 20150226_225951

PockelsCellsTransmission test

To check the proper behaviour after 10 hours of rest. Scan between 0.1 and 2.0us in 7 steps with IS:

TN Sup Plot
''20150226_230317, '20150226_230557' B=82, Y=122, R=97 20150226_230317_pcTransmission.jpg

Subapertures redefinition

Yesterday the loop was diverging for high order gain > 0.1. Most of the time the failure was triggered by the central actuators around the central obstruction. We try to redefine aq proper subaperture set avoiding the certer cross subapertures around the obstruction.

On Yellow we need to remove: 251, 265, 266, 267, 281

On Blue we need to remove: 74, 88, 89, 90, 104

On Red we need to remove: 428, 442, 443, 444, 458

Since we want to add 2 edge subapertures (527, 528) on the Red one we try to start from scratch.

We need also to remove the top row of the Blue pupil.

Subap CoG: 20150226_233800 -> BAD top row in Blue pupil is WRONG!

Subap TCoG48: 20150226_234900 -> BAD top row in Blue pupil is WRONG!

Reference Pupil and Piezo position for IM

Pupil PM X PM Y PI X PI Y
Blue 755 514 -1.5 -0.3
Yellow 288 723 -2.6 2.4
Red 455 663 -0.3 -2.3
PUPIL STABILIZATION TARGET: 20150227_002749

LGSW to ASM interaction matrix

We calibrate directly 150modes closing the loop with the old reconstructors.

TCoG48

IM Shift REC Notes
20150227_003817 4 20150227_004500 TT removed
Then we try to record another IM but instead of subtracting the TT component from the higher order modes we close the AO loop on the LGSW using a sensible gain on the TT terms:

IM Shift REC Notes
20150227_005122 4 20150227_005800 TT CL on LGSW, TTg= 0.5
20150227_010315 4 20150227_011900 TT CL on LGSW, TTg= 0.3
Now we combine the RECs:

Combination Tag Subaps LGSW REC TT REC Combined REC
TCoG48 only, TTg=0.3 20141126_212700 20150227_011900 - 20150227_012100
TCoG48 only, TT RM 20141126_212700 20150227_004500 - 20150227_012200

RECs test

We try to close the loop using a disturb = 1.2". We try to make ASM oscillates to check the maximum gain applicable (g=0.9 in case of TT RM). We compare the TT RM and TT g=0.3 RECs.

TN AO Gain Res. WF [nm rms] AO rec JCL PCL Notes Plot
20150227_012955 -0.8,-0.8,-0.8 726.593 20150227_012200 F F TT RM 20150227_012955-20150227_013748_modalplot.png
20150227_013119 -0.7,-0.7,-0.6 122.333 20150227_012200 F F
20150227_013201 -0.5,-0.5,-0.5 156.955 20150227_012200 F F
20150227_013705 -0.8,-0.8,-0.8 130.047 20150227_012100 F F TT g=0.3
20150227_013748 -0.5,-0.5,-0.5 149.623 20150227_012100 F F
The 2 RECs seems to have similar performaces both at g=0.5 and g=0.8...

200 modes control test

IM Shift REC Notes
20150227_014749 4 20150227_015000 200 modes, TT RM, TCoG48

Combination Tag Subaps LGSW REC TT REC Combined REC
TCoG48 only, 200 modes 20141126_212700 20150227_015000 - 20150227_015100

TN AO Gain Res. WF [nm rms] AO rec JCL PCL Notes Plot
20150227_015350 -0.5,-0.5,-0.3 257.651 20150227_015100 F F 200 modes controlled 20150227_015350-20150227_015350_modalplot.png

BUG in REC generation & new RECs

It is caused by indexing problems, discovered multiplying the REC by the IM (TT removed). Figures below show the result of the matrices multiplication in case of the old REC (20141202_125200) and of the newly generated REC (20150227_023600):

REC w/ indexing BUG REC w/o indexing BUG
rec20141202_125200_im_logscale.png rec20150227_023600_im_logscale.png

IM Shift REC Notes
20150227_003817 4 20150227_023500 TT RM, TCoG48
20150227_032211 4 ??? TCoG48, 200 modes

Combination Tag Subaps LGSW REC TT REC Combined REC
TCoG48 only 20141126_212700 20150227_023500 - 20150227_023600
TCoG48 BIN2 20141126_212700 20150227_023500 20130916_221257 20150227_023800
TCoG48 BIN4 20141126_212700 20150227_023500 20130917_214654 20150227_023900
TCoG48 QC 20141126_212700 20150227_023500 20150223_095600 20150227_024000

TN AO Gain Res. WF [nm rms] AO rec JCL PCL Notes Plot
20150227_031321 -0.4,-0.5,-0.5 169.871 20150227_023600 F F Disturb=0.6" 20150227_031321-20150227_032211_modalplot.png
20150227_031404 -0.4,-0.5,-0.5 182.69 20150227_023600 F F
20150227_032211 -0.4,-0.5,-0.5 157.151 20150227_023600 F F Disturb=1.2"

NightTime activity

Hip 40075

22:00: we close the loop on Hip 40075, m=13 on PYR, FW1=50-50. LGSW + QC loop. We have problems in keeping the loop closed: the edges of the pupils disappear and the shell RIPs... Other times we have the whole central stripe disappearing...

TN AO Gain Res. WF [nm rms] AO rec JCL PCL Notes
20150227_050142 0.7,-0.3, -0 196.961 20150227_024000 F F 20150227_050142-20150227_054522_modalplot.png
20150227_051912 0, 0, 0 726.786 20141127_222222 F F
20150227_053852 0.1,-0.1, 0 656.622 20141127_222222 F F
20150227_053937 0.1,-0.1, 0 582.591 20141127_222222 F F OL
20150227_054015 0.1,-0.1, 0 614.525 20141127_222222 F F OL
20150227_054242 0.1,-0.1, 0 724.834 20141127_222222 F F OL
20150227_054417 0, 0, 0 765.914 20141127_222222 F F OL
20150227_054522 0.2,-0.15, -0 621.198 20141127_222222 F F

Hip 67995

3:00: after having optimized the LGS spots the loop is more stable... Seeing is about 1.2". QC is running at 1kHz, 1300 counts. TS is on, PYR BIN4, 1kHz, 70 counts/subap, m=12.6, FW1=50-50.

After the gain is stable to 0.6,0.4,0.2 we take LUCI images range from 0008 to 0017 are in H band (3s), images in the range 0019 - 0028 are in FeII band (3s).

Focus -150k
Corr 750
Rot 6500

TN AO Gain Res. WF [nm rms] AO rec JCL PCL Notes
20150227_103733 -0.3,-0.4,-0.2 2302.66 20150227_023600 F F LGSW + QC 20150227_103733-20150227_105656_modalplot.png
20150227_103758 -0.3,-0.4,-0.2 2236.21 20150227_023600 F F
20150227_103838 -0.3,-0.4,-0.2 2369.19 20150227_023600 F F
20150227_104404 0.5,-0.4,-0.2 159.709 20150227_024000 T F
20150227_104427 0.5,-0.4,-0.2 170.583 20150227_024000 T F
20150227_104639 0.6,-0.4,-0.2 167.03 20150227_024000 T F
20150227_104704 0.6,-0.4,-0.2 158.636 20150227_024000 T F
20150227_105255 0.6,-0.4,-0.2 179.194 20150227_024000 T F
20150227_105345 0.6,-0.4,-0.2 162.144 20150227_024000 T F
20150227_105656 0.6,-0.5,-0.3 635.053 20150227_024000 T F

VibrationSystemDx test

We close the TT loop on a bright star, the pupil mirror starts to oscillate in different directions, we record the proper LGSW signalswith snapshots. We have also a set of LUCI images 0029- 0038 when just the TT loop was closed (field stabilization).

TN AO Gain Res. WF [nm rms] AO rec JCL PCL Notes
20150227_110622 0.5, -0, -0 467.898 20150227_024000 F F
20150227_110744 0.5, -0, -0 356.964 20150227_024000 F F
20150227_110931 0.5, -0, -0 577.949 20150227_024000 F F
20150227_110948 0.5, -0, -0 451.557 20150227_024000 F F
20150227_111032 0.5, -0, -0 345.892 20150227_024000 F F
20150227_111127 0.5, -0, -0 474.255 20150227_024000 F F
20150227_111215 0.5, -0, -0 387.213 20150227_024000 F F
20150227_111321 0.5, -0, -0 308.894 20150227_024000 F F
20150227_111401 0.5, -0, -0 434.998 20150227_024000 F F
20150227_111431 0.5, -0, -0 473.871 20150227_024000 F F
20150227_111436 0.5, -0, -0 404.933 20150227_024000 F F
20150227_111509 0.5, -0, -0 316.451 20150227_024000 F F
20150227_111525 0.5, -0, -0 357.229 20150227_024000 F F
20150227_111549 0.5, -0, -0 370.301 20150227_024000 T F
20150227_111619 0.5, -0, -0 331.774 20150227_024000 T F
20150227_111632 0.5, -0, -0 389.297 20150227_024000 T F
20150227_111701 0.5, -0, -0 333.413 20150227_024000 T F
20150227_111731 0.5, -0, -0 365.498 20150227_024000 T F
20150227_111802 0.5, -0, -0 401.62 20150227_024000 T F
20150227_110622-20150227_111802_modalplot.png

Hip 67995 (field stabilization mode)

Here we are running just the TT loop and we are taking images of Hip 67995 with LUCI in different conditions

LUCI TN Notes
0040 - 0049 QC g=2.0
0050 - 0059 QC g=2.0, loop is stable
0060 - 0069 just guiding, g=0.001

TN AO Gain Res. WF [nm rms] AO rec JCL PCL Notes
20150227_113327 2, -0, -0 338.695 20150227_024000 F F
We reoptimize the spot shape

Focus -155k
Corr 1000
Rot 6500
We close againwith the full system:

TN AO Gain Res. WF [nm rms] AO rec JCL PCL Notes Plot
20150227_122135 1.5,-0.5,-0.3 133.603 20150227_024000 T F LUCI TN: 0072- 0081 20150227_122135-20150227_122354_modalplot.png
20150227_122354 1.5,-0.5,-0.3 137.673 20150227_024000 T F LUCI TN: 0082
20150227_123008 0.005, -0, -0 443.658 20150227_024000 T F LUCI TN: 0083 - 0092 20150227_123008-20150227_123212_modalplot.png
20150227_123026 0.005, -0, -0 400.303 20150227_024000 T F
20150227_123046 0.005, -0, -0 321.669 20150227_024000 T F
20150227_123048 0.005, -0, -0 310.67 20150227_024000 T F
20150227_123113 0.005, -0, -0 362.827 20150227_024000 T F
20150227_123134 0.005, -0, -0 368.084 20150227_024000 T F
20150227_123212 0.005, -0, -0 345.137 20150227_024000 T F
20150227_123530 3,-0.6,-0.5 164.123 20150227_024000 T F LUCI TN: 0093 - 0102 20150227_123530-20150227_123722_modalplot.png
20150227_123544 3,-0.6,-0.5 170.499 20150227_024000 T F
20150227_123610 3,-0.6,-0.5 244.92 20150227_024000 T F
20150227_123614 3,-0.6,-0.5 269.847 20150227_024000 T F
20150227_123623 3,-0.6,-0.5 271.178 20150227_024000 T F
20150227_123643 3,-0.6,-0.5 392.069 20150227_024000 T F
20150227_123704 2,-0.6,-0.5 145.137 20150227_024000 T F
20150227_123722 2,-0.6,-0.5 144.336 20150227_024000 T F

LUCI TN Notes
0072 - 0081 FeII, 3s
0082 FeII, 10s
0083 - 0092 FeII, 3s
0093 - 0102 FeII, 3s

NCG 5272

6:00: we point to NCG 5272.

LGSW + QC + TS. m=13 on PYR, FW1=50-50, TS running at 100Hz, 400 counts /subap. QC at 1kHz, 900 counts

TN AO Gain Res. WF [nm rms] AO rec JCL PCL Notes Plot
20150227_130159 1.5,-0.6,-0.3 163.735 20150227_024000 T F 20150227_130159-20150227_130238_modalplot.png
20150227_130238 1.5,-0.6,-0.5 174.089 20150227_024000 T F

NCG 5272 (dithering)

We manually move the FLAO stages to dither the NCG5272 image

TN AO Gain Res. WF [nm rms] AO rec JCL PCL Notes Plot
20150227_130951 1.5,-0.6,-0.5 168.721 20150227_024000 T F [-74;-50], LUCI TN: 0155 - 0162, K band 20150227_130951-20150227_130951_modalplot.png
20150227_131423 1.5,-0.6,-0.5 134.697 20150227_024000 T F [-74;-51], LUCI TN: 0163 - 0167
20150227_131649 1.5,-0.6,-0.5 165.801 20150227_024000 T F [-75;-51], LUCI TN: 0168 - 0172
20150227_131941 1.5,-0.6,-0.5 163.367 20150227_024000 T F [-76;-51], LUCI TN: 0173 - 0177
20150227_132230 1.5,-0.6,-0.5 139.207 20150227_024000 T F [-77;-51], LUCI TN: 0178 - 0182
20150227_132413 1.5,-0.6,-0.5 136.086 20150227_024000 T F [-77;-50], LUCI TN: 0183 - 0187
20150227_132601 1.5,-0.6,-0.5 120.81 20150227_024000 T F [-77;-49], LUCI TN: 0188 - 0192, J band

-- MarcoBonaglia - 05 Mar 2015
Topic attachments
I Attachment Action Size Date Who Comment
20150226_230317_pcTransmission.jpgjpg 20150226_230317_pcTransmission.jpg manage 41 K 05 Mar 2015 - 11:30 MarcoBonaglia  
20150227_012955-20150227_013748_modalplot.pngpng 20150227_012955-20150227_013748_modalplot.png manage 56 K 05 Mar 2015 - 14:34 MarcoBonaglia  
20150227_015350-20150227_015350_modalplot.pngpng 20150227_015350-20150227_015350_modalplot.png manage 27 K 05 Mar 2015 - 14:39 MarcoBonaglia  
20150227_031321-20150227_031404_modalplot.pngpng 20150227_031321-20150227_031404_modalplot.png manage 38 K 05 Mar 2015 - 14:59 MarcoBonaglia  
20150227_031321-20150227_032211_modalplot.pngpng 20150227_031321-20150227_032211_modalplot.png manage 51 K 05 Mar 2015 - 15:02 MarcoBonaglia  
20150227_050142-20150227_054522_modalplot.pngpng 20150227_050142-20150227_054522_modalplot.png manage 75 K 05 Mar 2015 - 15:07 MarcoBonaglia  
20150227_103733-20150227_105656_modalplot.pngpng 20150227_103733-20150227_105656_modalplot.png manage 69 K 05 Mar 2015 - 15:14 MarcoBonaglia  
20150227_110622-20150227_111802_modalplot.pngpng 20150227_110622-20150227_111802_modalplot.png manage 116 K 05 Mar 2015 - 15:17 MarcoBonaglia  
20150227_122135-20150227_122354_modalplot.pngpng 20150227_122135-20150227_122354_modalplot.png manage 58 K 05 Mar 2015 - 15:32 MarcoBonaglia  
20150227_123008-20150227_123212_modalplot.pngpng 20150227_123008-20150227_123212_modalplot.png manage 76 K 05 Mar 2015 - 15:39 MarcoBonaglia  
20150227_123530-20150227_123722_modalplot.pngpng 20150227_123530-20150227_123722_modalplot.png manage 77 K 05 Mar 2015 - 15:41 MarcoBonaglia  
20150227_130159-20150227_130238_modalplot.pngpng 20150227_130159-20150227_130238_modalplot.png manage 60 K 05 Mar 2015 - 15:55 MarcoBonaglia  
20150227_130951-20150227_130951_modalplot.pngpng 20150227_130951-20150227_130951_modalplot.png manage 79 K 05 Mar 2015 - 15:54 MarcoBonaglia  
20150227_130951-20150227_132601_modalplot.pngpng 20150227_130951-20150227_132601_modalplot.png manage 85 K 05 Mar 2015 - 15:52 MarcoBonaglia  
rec20141202_125200_im_logscale.pngpng rec20141202_125200_im_logscale.png manage 73 K 05 Mar 2015 - 14:44 MarcoBonaglia  
rec20150227_023600_im_logscale.pngpng rec20150227_023600_im_logscale.png manage 56 K 05 Mar 2015 - 14:44 MarcoBonaglia  
Topic revision: r3 - 18 Mar 2015, TommasoMazzoni
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback